# A Bayesian hierarchical mixture cure modelling framework to utilize multiple survival datasets for long-term survivorship estimates: A case study from previously untreated metastatic melanoma

### v1

@@ Abstract
Time to an event of interest over a lifetime is a central measure of the clinical benefit of an intervention used in a health technology assessment (HTA). Within the same trial multiple end-points may also be considered. For example, overall and progression-free survival time for different drugs in oncology studies. A common challenge is when an intervention is only effective for some proportion of the population who are not clinically identifiable. Therefore, latent group membership as well as separate survival models for groups identified need to be estimated. However, follow-up in trials may be relatively short leading to substantial censoring. We present a general Bayesian hierarchical framework that can handle this complexity by exploiting the similarity of cure fractions between end-points; accounting for the correlation between them and improving the extrapolation beyond the observed data. Assuming exchangeability between cure fractions facilitates the borrowing of information between end-points. We show the benefits of using our approach with a motivating example, the CheckMate 067 phase 3 trial consisting of patients with metastatic melanoma treated with first line therapy.
@@ 1 Introduction
Interventions that impact the time to an event of interest, such as disease progression or death in a cancer trial, form a high proportion of appraisals by Health Technology Assessment (HTA) agencies, such as the National Institute for Health and Care Excellence (NICE), in England [1]. For instance, a new intervention may be expected to increase the amount of time until a patient experiences disease progression or death from any cause. For this reason, “survival data” are instrumental to modelling in HTA [2, 3].
@@
In order to quantify accurately the health and economic benefits of a new intervention using such survival data, it is necessary to estimate the mean survival time, ie, the long-term effects of a given intervention. This is in contrast to general survival curve summaries, such as the median time, which are typically used in standard “biostatistical” analyses. This idiosyncrasy has important implications because, in order to calculate the mean survival time we require the full survival curve, ie, over a patient’s lifetime, but available data (e.g. data-cuts from a randomised trial) almost inevitably only cover a limited time frame and are subject to a high degree of censoring. Thus, it is necessary to extrapolate the observed survival curve to a time horizon that is typically much longer than the one considered in the experimental study. Consequently, unlike “standard” time-to-event analyses that are often based on semi-parametric models (e.g. Cox regression), HTA modelling is based on a fully parametric approach to the survival analysis, as recommended by a highly influential NICE Decision Support Unit (DSU) Technical Support Document (TSD) [1].
@@
Another important feature of survival data in the field of cancer research has arisen in the past decade due to the development of potentially highly successful immuno-oncology treatments. These aim to stimulate the body’s immune system to recognise and kill cancer cells [4]. Several types of immunotherapy can be used to treat melanoma including immune checkpoint inhibitors, interleukin-2 (IL-2), and oncolytic virus therapy. In particular, inhibitory immune checkpoint blockade combination therapies have shown significant success in promoting immune responses against cancer and can result in tumour regression in many patients [5]. These advances have lead to cancer patients with improved survival end-points and more long-term survival (LTS) (although complete responders, CR, still have worse outcomes than the general population). In previous immuno-oncologic studies for melanoma therapies, such as those evaluating ipilimumab and nivolumab, results have indicated that survival curves “plateau” for a considerable proportion of LTS patients in the time period covered by survival curve [6, 7]. That is, the underlying survival curve converges to a probability (substantially) greater than zero and does not appear to decrease further. Note that in practice, in the tail of the Kaplan-Meier plateauing may also be due to sampling issues, such as small sample size or missing data, and so there will be uncertainty about the true proportion of LTS. When a survival curve exhibits this plateau behaviour, it is common to adopt a mixture cure model (MCM) approach to describe the generation of the underlying data. A MCM considers a population as a mixture of two groups: the cured and not cured. In our case, the cured group contains the LTS patients. In many MCM analyses the survival associated with the uncured fraction of the population is represented by a Cox proportional hazard model.
@@
However, the survival plateaus indicating a proportion of LTS may not have been reached at the time of a data-cut (including the planned end of the study). In order to obtain complete survival curves to use in an HTA evaluation we need to extrapolate the survival curves to the sustained plateaus. Ideally, these extrapolations should be consistent between data-cuts [8]. The question of how to principally and usefully do this in situations where trials include multiple treatment arms and end-points is the subject of this paper. The fact that we consider multiple end-points within a trial in the same analysis means that information about the plateau from data for one end-point can potentially help inform the estimation of the plateau for another end-point. This can enable consistency between plateau estimates and maximise the utility of all available data.
@@
Within a frequentist paradigm, an example of related work is a multi-level modelling approach in mixture cure modelling using random effects to model multilevel clustering structure in the linear predictors in both hazard function and cured probability parts [9]. Correlation between random effects in the uncured survival and cure fraction has been investigated using a bivariate Normal distribution [10] [11]. Instead, we will impose dependency with hyperparameters. To our knowledge there has been no Bayesian fully-parametric mixture cure model with a multi-level modelling structure for end-points.
@@ 2 Motivating example
Our motivating example concerns a long-term study of melanoma therapies. The study has been described in detail elsewhere, so we give a brief overview of the data here (see [6, 7, 12] for more details). The CheckMate 067 trial is a phase 3 randomised, double-blind study conducted on eligible patients aged 18 years or older with previously untreated, unresectable stage III or IV melanoma. Nivolumab alone or the combination of nivolumab plus ipilimumab was compared with ipilimumab alone in patients with metastatic melanoma. They were randomly assigned in a 1:1:1 ratio and stratified by anti–programmed death-1 ligand 1 (PD-L1) expression status, whether there is a mutation in the BRAF gene that makes it work incorrectly (BRAF mutation), and AJCC metastasis stage. The total baseline sample consists of $n=945$ subjects divided across three treatment arms i) ipilimumab monotherapy with 3 mg/kg IV once every 3 weeks (Q3W) for a total of 4 doses; ii) nivolumab monotherapy with 3 mg/kg intravenous (IV) once every 2 weeks (Q2W); iii) combined nivolumab with ipilimumab with 1 mg/kg IV and 3 mg/kg IV Q3W, respectively, for 4 doses followed by nivolumab 3 mg/kg IV Q2W. The co-primary end-points were progression-free survival (PFS; ie, the time from randomisation until the first of disease progression or death) and overall survival (OS; ie, the time from randomisation to time of death). The data source for the analysis consists of patient-level data of times to PFS and OS with covariates sex, age at start of trial, and country of centre.
@@
The reader is encouraged to refer to previously published analyses for the CheckMate 067 trial results, including the 5-year database lock (DBL) publication[7]. In summary, for PFS the median (95% CI) times to event for each treatment arm were 2.86 months (2.79-3.29) for ipilimumab, 6.93 months (5.32-10.41) for nivolumab, and 11.50 months (9.26-20.80) for the combination treatment of nivolumab and ipilimumab. For OS the median (95% CI) times to event for each treatment arm were 20.0 months (17.22-25.6) for ipilimumab, 36.9 months (31.24-60.9) for nivolumab, and not observed for the combination treatment. Figure 1 shows the Kaplan-Meier estimates of the PFS and OS survival curves for the complete CheckMate 067 phase 3 trial data. Both sets of curves, for PFS and OS, appear to approach a positive limiting probability; the PFS curves more clearly so. Early published results for outcomes at 7.5 years follow-up continue to demonstrate the durability of responses with combined nivolumab and ipilimumab, and an ongoing survival plateau [13].
@@
A standard frequentist analysis has been carried-out on these data already with separate mixture cure models for each treatment and OS or PFS [14]. It was estimated that for OS, ranges of estimated proportions of LTSs were 16%–26% for ipilimumab, 38%–46% for nivolumab, and 49%–54% for combined treatment across all modelled distributions. Similarly, for PFS, ranges of the proportions of LTSs were 9%–13% for ipilimumab, 29%–33% for nivolumab, and 38%–40% for combined treatment.
@@ 3 Modelling framework
Consider a study in which $n$ individuals are observed, and the (non-negative) time to either progression ($T^{prog}_{i}$) or death ($T^{death}_{i}$) are recorded for each individual $i=1,\ldots,n$. Denote the event time for individual $i$ by $T_{i}$. In our motivating example, we would have a separate set of $T_{i}$ for either end-point, PFS or OS. Thus, PFS times are $T_{i}=\min(T^{prog}_{i},T^{death}_{i})$, and OS times are $T_{i}=T^{death}_{i}$. Event times may be censored, and therefore we consider a censoring time $C_{i}$ and an event indicator $\delta_{i}=\mathbb{I}(T_{i}<C_{i})$ — so that $\delta_{i}=1$ if the end-point is fully observed for individual $i$ and 0 if it is censored. The observed survival time is then indicated as $t_{i}=\min(T_{i},C_{i})$.
Typically, a study will also measure some individual-level covariates, which we assume are collected in a vector $\bm{x}_{i}=(x_{1i},\ldots,x_{Pi})^{\top}$. These may include the individual’s age, sex, comorbidities, and usually, a treatment arm indicator. The full data can then be represented as $\mathcal{D}_{i}=(t_{i},\delta_{i},\bm{x}_{i})$. In our example, the covariates are the same between PFS and OS.
We model the observed times using a parametric sampling distribution $t_{i}\sim p(t_{i}\mid\bm{\theta})$ as a function of a set of model parameters $\bm{\theta}=(\bm{\alpha}(\bm{x}),\mu(\bm{x}))$. In this notation, the location parameter $\mu(x)$ indicates the mean or the scale of the probability distribution; and a (set of) ancillary parameters $\bm{\alpha}(x)$ describe the shape or variance of the distribution. Whilst it is possible for both $\mu$ and $\bm{\alpha}$ to explicitly depend on the covariates $\bm{x}$, usually the formulation is simplified to assume that these only affect directly the location parameter. Since $t>0$, we typically use a generalised linear formulation. So, for an individual $i$
$g(\mu)=\beta^{\mu}_{0}+\sum_{p=1}^{P}\beta^{\mu}_{p}x_{pi}\;[+\ldots],$
to model the rate or location parameter, where $\beta^{\mu}_{0}$ represents an intercept, and $\beta^{\mu}_{p}$ are the coefficients for an individual’s measurements. The [+]delimited-[][$\ldots$][ + … ] term indicates additional components for the linear predictor (e.g. structured random effects). The function $g(\cdot)$ is typically the logarithm.
@@ 3.1 The standard mixture cure model
Suppose that we observe a sustained plateau in the survival time data or have some other reason to believe that there is a long-term survivor group in our sample. We will now present the standard mixture cure model (sometimes called a long-term survival model) as the basis of our proposed extension. Denote the probability of being cured by the cure fraction, $\pi$. The term ”cure” here refers to the type of statistical model and does not literally mean that they are cured of disease. In our case, cured means that a patient has responded to treatment and is a LTS. Because the cure models explicitly split patients in to two groups (cured or not) then the full set of model parameters consists of two separate groups. Define $\bm{\omega}=(\bm{\theta}^{b},\bm{\theta}^{u})$, where superscript $b$ denotes the background model and $u$ denotes the uncured model. In the standard cure model, cured individuals are usually assumed to have no chance of experiencing the event of interest in the period under consideration, simplifying the model. There are several types of cure models and a rich and growing body of research in this area. (see [15] for a comparison and guidance with application to oncology.)
@@
where $S(t)=1\!-\!\int_{0}^{t}p(s\mid\theta)\,\text{d}s$ denotes survival at time $t$, $S_{b}(t\mid\bm{\theta}^{b},\bm{x})$ is a function of the background mortality at time $t$ conditional on covariates $\bm{x}$, and $S_{u}(t\mid\bm{\theta}^{u},\bm{x})$ is a function of the (excess) mortality due to cancer at time $t$, conditional on covariates $\bm{x}$. The formulation in equation (1) can be thought of as comprising of two submodels: i) an incidence model for the cure fraction $\pi$ (the general term from epidemiology is the rate or probability of a disease); and ii) a latency model for the survival function of the uncured population $S_{u}$. This is ’hidden’ since it cannot be directly observed from the data.
@@
$\mbox{logit}(\pi_{k})=\beta^{\pi}_{k}\;[+\ldots],$ (2)
where $\beta^{\pi}_{k}$ are the regression coefficients quantifying the impact of treatment $k$. The [+]delimited-[][$\ldots$][ + … ] term could include a frailty term in the model. The treatment index for an individual is included in the covariate vector $\bm{x}$ so that $\pi$ is replaced by $\pi(\bm{x})$ in equation (1). Note that although we are including all treatments arms in the same model we are not assuming any relationship between them. In particular, we do not suppose proportional hazards. This allows greater flexibility for the cases when proportional hazards is violated.

### v2

# N Green et al

@@ Abstract
[Summary] Time to an event of interest over a lifetime is a central measure of the clinical benefit of an intervention used in a health technology assessment (HTA). Within the same trial, multiple end-points may also be considered. For example, overall and progression-free survival time for different drugs in oncology studies. A common challenge is when an intervention is only effective for some proportion of the population who are not clinically identifiable. Therefore, latent group membership as well as separate survival models for identified groups need to be estimated. However, follow-up in trials may be relatively short leading to substantial censoring. We present a general Bayesian hierarchical framework that can handle this complexity by exploiting the similarity of cure fractions between end-points; accounting for the correlation between them and improving the extrapolation beyond the observed data. Assuming exchangeability between cure fractions facilitates the borrowing of information between end-points. We undertake a comprehensive simulation study to evaluate the model performance under different scenarios. We also show the benefits of using our approach with a motivating example, the CheckMate 067 phase 3 trial consisting of patients with metastatic melanoma treated with first line therapy.
[…]
keywords:
Survival analysis, multi-level model, extrapolation, oncology
articletype: Article Type00footnotetext: Abbreviations: MCM, mixture cure model
@@ 1 Introduction
Interventions that impact the time to an event of interest, such as disease progression or death in a cancer trial, form a high proportion of appraisals by Health Technology Assessment (HTA) agencies, such as the National Institute for Health and Care Excellence (NICE), in England 1. For instance, a new intervention may be expected to increase the amount of time until a patient experiences disease progression or death from any cause. For this reason, “survival data” are instrumental to modelling in HTA 2, 3.
@@
In order to quantify accurately the health and economic benefits of a new intervention using such survival data, it is necessary to estimate the mean survival time, ie, the long-term effects of a given intervention. This is in contrast to general survival curve summaries, such as the median time, which are typically used in standard “biostatistical” analyses. This idiosyncrasy has important implications because, in order to calculate the mean survival time we require the full survival curve, ie, over a patient’s lifetime, but available data (e.g. data-cuts from a randomised trial) almost inevitably only cover a limited time frame and are subject to a high degree of censoring. Thus, it is necessary to extrapolate the observed survival curve to a time horizon that is typically much longer than the one considered in the experimental study. Consequently, unlike “standard” time-to-event analyses that are often based on semi-parametric models (e.g. Cox regression), HTA modelling is based on a fully parametric approach to the survival analysis, as recommended by a highly influential NICE Decision Support Unit (DSU) Technical Support Document (TSD) 1.
@@
Another important feature of survival data in the field of cancer research has arisen in the past decade due to the development of potentially highly successful immuno-oncology treatments. These aim to stimulate the body’s immune system to recognise and kill cancer cells 4. Several types of immunotherapy can be used to treat melanoma including immune checkpoint inhibitors, interleukin-2 (IL-2), and oncolytic virus therapy. In particular, inhibitory immune checkpoint blockade combination therapies have shown significant success in promoting immune responses against cancer and can result in tumour regression in many patients 5. These advances have lead to cancer patients with improved survival end-points and more long-term survival (LTS) (although complete responders, CR, still have worse outcomes than the general population). In previous immuno-oncologic studies for melanoma therapies, such as those evaluating ipilimumab and nivolumab, results have indicated that survival curves “plateau” for a considerable proportion of LTS patients in the time period covered by survival curve 6, 7. That is, the underlying survival curve converges to a probability (substantially) greater than zero and does not appear to decrease further. Note that in practice, in the tail of the Kaplan-Meier plateauing may also be due to sampling issues, such as small sample size or missing data, and so there will be uncertainty about the true proportion of LTS. When a survival curve exhibits this plateau behaviour, it is common to adopt a mixture cure model (MCM) approach to describe the generation of the underlying data. A MCM considers a population as a mixture of two groups: the cured and not cured. In our case, the cured group contains the LTS patients. In many MCM analyses the survival associated with the uncured fraction of the population is represented by a Cox proportional hazard model. Alongside this consideration, it is also important to account for the additional risk experienced by the uncured patients due to the cancer, beyond the background mortality risk of the general population (the cured). Excess hazard models provide a framework to separate the baseline mortality (background hazard) from the additional risk associated with the disease. This can be especially important when dealing with survival data where patients may not all have the same risk of death.
@@
However, the survival plateaus indicating a proportion of LTS may not have been reached at the time of a data-cut (including the planned end of the study). In order to obtain complete survival curves to use in an HTA evaluation we need to extrapolate the survival curves to the sustained plateaus. Ideally, these extrapolations should be consistent between data-cuts 8. The question of how to principally and usefully do this in situations where trials include multiple treatment arms and end-points is the subject of this paper. The fact that we consider multiple end-points within a trial in the same analysis means that information about the plateau from data for one end-point can potentially help inform the estimation of the plateau for another end-point. This can enable consistency between plateau estimates and maximise the utility of all available data.
@@
Within a frequentist paradigm, an example of related work is a multi-level modelling approach in mixture cure modelling using random effects to model multilevel clustering structure in the linear predictors in both the hazard function and cured probability parts 9. Correlation between random effects in the uncured survival and cure fraction has been investigated using a bivariate Normal distribution 10 11. Instead, we will impose dependency with hyperparameters. To our knowledge there has been no Bayesian fully-parametric mixture cure model with a multi-level modelling structure for end-points.
@@ 2 Motivating example
Our motivating example concerns a long-term study of melanoma therapies. The study has been described in detail elsewhere, so we give a brief overview of the data here (see 6, 7, 12 for more details). The CheckMate 067 trial is a phase 3 randomised, double-blind study conducted on eligible patients aged 18 years or older with previously untreated, unresectable stage III or IV melanoma. Nivolumab alone or the combination of nivolumab plus ipilimumab was compared with ipilimumab alone in patients with metastatic melanoma. They were randomly assigned in a 1:1:1 ratio and stratified by anti–programmed death-1 ligand 1 (PD-L1) expression status, whether there is a mutation in the BRAF gene that makes it work incorrectly (BRAF mutation), and AJCC metastasis stage. The total baseline sample consists of $n=945$ subjects divided across three treatment arms i) ipilimumab monotherapy with 3 mg/kg IV once every 3 weeks (Q3W) for a total of 4 doses; ii) nivolumab monotherapy with 3 mg/kg intravenous (IV) once every 2 weeks (Q2W); iii) combined nivolumab with ipilimumab with 1 mg/kg IV and 3 mg/kg IV Q3W, respectively, for 4 doses followed by nivolumab 3 mg/kg IV Q2W. The co-primary end-points were progression-free survival (PFS; ie, the time from randomisation until the first of disease progression or death) and overall survival (OS; ie, the time from randomisation to time of death). Disease progression was defined as evaluated tumour response using criteria from RECIST (Response Evaluation Criteria in Solid Tumors), version 1.1. This includes a significant increase in size, new lesions or worsening of existing lesions. The data source for the analysis consists of patient-level data of times to PFS and OS with covariates sex, age at the start of the trial, and country of centre.
@@
The reader is encouraged to refer to previously published analyses for the CheckMate 067 trial results, including the 5-year database lock (DBL) publication7. In summary, for PFS the median (95% CI) times to event for each treatment arm were 2.86 months (2.79-3.29) for ipilimumab, 6.93 months (5.32-10.41) for nivolumab, and 11.50 months (9.26-20.80) for the combination treatment of nivolumab and ipilimumab. For OS the median (95% CI) times to event for each treatment arm were 20.0 months (17.22-25.6) for ipilimumab, 36.9 months (31.24-60.9) for nivolumab, and not observed for the combination treatment. Figure 1 shows the Kaplan-Meier estimates of the PFS and OS survival curves for the complete data from the CheckMate 067 phase 3 trial. Both sets of curves, for PFS and OS, appear to approach a positive limiting probability; the PFS curves more clearly so. Early published results for outcomes at 7.5 years follow-up continue to demonstrate the durability of responses with combined nivolumab and ipilimumab, and an ongoing survival plateau 13.
@@
A standard frequentist analysis has been carried-out on these data already with separate mixture cure models for each treatment and OS or PFS 14. It was estimated that for OS, ranges of estimated proportions of LTSs were 16%–26% for ipilimumab, 38%–46% for nivolumab, and 49%–54% for combined treatment across all modelled distributions. Similarly, for PFS, ranges of the proportions of LTSs were 9%–13% for ipilimumab, 29%–33% for nivolumab, and 38%–40% for combined treatment.
@@ 3 Modelling framework
Consider a study in which $n$ individuals are observed, and the (non-negative) time to either progression ($T^{prog}_{i}$) or death ($T^{death}_{i}$) is recorded for each individual $i=1,\ldots,n$. Denote the event time for individual $i$ by $T_{i}$. In our motivating example, we would have a separate set of $T_{i}$ for either end-point, PFS or OS. Thus, PFS times are $T_{i}=\min(T^{prog}_{i},T^{death}_{i})$, and OS times are $T_{i}=T^{death}_{i}$. Event times may be censored, and therefore we consider a censoring time $C_{i}$ and an event indicator $\delta_{i}=\mathbb{I}(T_{i}<C_{i})$ — so that $\delta_{i}=1$ if the end-point is fully observed for individual $i$ and 0 if it is censored. The observed survival time is then indicated as $t_{i}=\min(T_{i},C_{i})$.
Typically, a study will also measure some individual-level covariates, which we assume are collected in a vector $\bm{x}_{i}=(x_{1i},\ldots,x_{Pi})^{\top}$. These may include the individual’s age, sex, comorbidities, and usually, a treatment arm indicator. The complete data can then be represented as $\mathcal{D}_{i}=(t_{i},\delta_{i},\bm{x}_{i})$. In our example, the covariates are the same between PFS and OS.
We modelled the observed times using a parametric sampling distribution $t_{i}\sim p(t_{i}\mid\bm{\theta})$ as a function of a set of model parameters $\bm{\theta}=(\bm{\alpha}(\bm{x}),\mu(\bm{x}))$. In this notation, the location parameter $\mu(x)$ indicates the mean or scale of the probability distribution; and a (set of) ancillary parameters $\bm{\alpha}(x)$ describe the shape or variance of the distribution. Although it is possible for both $\mu$ and $\bm{\alpha}$ to explicitly depend on the covariates $\bm{x}$, usually the formulation is simplified to assume that these only directly affect the location parameter. Since $t>0$, we typically use a generalised linear formulation. So, for an individual $i$
$g(\mu)=\beta^{\mu}_{0}+\sum_{p=1}^{P}\beta^{\mu}_{p}x_{pi},$
to model the rate or location parameter, where $\beta^{\mu}_{0}$ represents an intercept, and $\beta^{\mu}_{p}$ are the coefficients for an individual’s measurements. The formula may also include additional components for the linear predictor (e.g. structured random effects). The function $g(\cdot)$ is typically the logarithm.
@@ 3.1 The standard mixture cure model
Suppose that we observe a sustained plateau in the survival time data or have some other reason to believe that there is a long-term survivor group in our sample. We will now present the standard mixture cure model (sometimes called a long-term survival model) as the basis of our proposed extension. Denote the probability of being cured by the cure fraction, $\pi$. The term "cure" here refers to the type of statistical model and does not literally mean that they are cured of disease. In our case, cured means that a patient has responded to treatment and is a LTS. Because the cure models explicitly split patients into two groups (cured or not) then the full set of model parameters consists of two separate groups. Define $\bm{\omega}=(\bm{\theta}^{b},\bm{\theta}^{u})$, where superscript $b$ denotes the background model and $u$ denotes the uncured model. In the standard cure model, cured individuals are usually assumed to have no chance of experiencing the event of interest in the period under consideration, simplifying the model. There are several types of cure models and a rich and growing body of research in this area. (see 15 for a comparison and guidance with application to oncology.)
@@
where $S(t)=1\!-\!\int_{0}^{t}p(s\mid\theta)\,\text{d}s$ denotes survival at time $t$, $S_{b}(t\mid\bm{\theta}^{b},\bm{x})$ is a function of the background mortality at time $t$ conditional on covariates $\bm{x}$, and $S_{u}(t\mid\bm{\theta}^{u},\bm{x})$ is a function of the (excess) mortality due to cancer at time $t$, conditional on covariates $\bm{x}$. The formulation in equation (1) can be thought of as comprising of two submodels: i) an incidence model for the cure fraction $\pi$ (the general term from epidemiology is the rate or probability of a disease); and ii) a latency model for the survival function of the uncured population $S_{u}$. This is ‘hidden’ since it cannot be directly observed from the data.
@@
$\mbox{logit}(\pi_{k})=\beta^{\pi}_{k},$ (2)
where $\beta^{\pi}_{k}$ are the regression coefficients quantifying the impact of treatment $k$. The formula could also contain additional terms, such as a frailty term. The treatment index for an individual is included in the covariate vector $\bm{x}$ so that $\pi$ is replaced by $\pi(\bm{x})$ in equation (1). Note that although we are including all treatment arms in the same model we are not assuming any relationship between them. In particular, we do not suppose proportional hazards. This allows greater flexibility for the cases when the proportional hazards assumption is violated.